OpenStudy (anonymous):

Select the choice below that could be the first step in solving the following system of equations using elimination: -6x − 7y = 12 3y − 4x − 1 = 0 Replace the right side of the first equation with the second equation. Multiply the second equation by 4. Multiply the second equation by -3. Write the second equation in terms of y.

OpenStudy (radar):

I would not use any of those choices. Are you sure you have copied this problem correctly?

OpenStudy (radar):

O.K. Slooooowwwwwww me, I see it now. You would multiply 2nd by a -3, then you would multiply the first equation by 2 so you could eliminate x.

OpenStudy (anonymous):

3y - 4x - 1 = 0 let me change this around -4x + 3y = 1 thats better....now we have : -6x - 7y = 12 -4x + 3y = 1 yes....I agree with radar.....you could multiply 2nd equation by -3 and then the first equation by 2, this would make the x's cancel out.

OpenStudy (radar):

A really odd way to ask this problem, I think it may contribute more to confusion than teaching, especially if you choose to solve the darn thing.!!